# The Complexity of Resilience Problems via Valued Constraint Satisfaction Problems

### v3

@@ 8.4. The Resilience Tractability Conjecture
Let $\mu$ be a union of connected conjunctive queries over the signature $\tau$, and let $\sigma\subseteq\tau$. If the structure ($(\{0,1\};\operatorname{OIT})$ has no pp-construction in $\Gamma:=\Gamma_{\mu,\sigma}$, then $\Gamma$ has a fractional polymorphism of arity $\ell\geq 2$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}(\Gamma)$ (and in this case, $\operatorname{VCSP}(\Gamma)$ is in P by Theorem 7.17).
The conjecture is intentionally only formulated for VCSPs that stem from resilience problems, because it is known to be false for the more general situation of VCSPs for valued structures $\Gamma$ that have the same automorphisms as a reduct of a finitely bounded homogeneous structure (Bodirsky, 2021) (Section 12.9.1; the counterexample is even a CSP). However, the structures ${\mathfrak{B}}_{\mu}$ from Theorem 8.14 that we need to formulate resilience problems as VCSPs are particularly well-behaved for the universal-algebraic approach and more specifically, for canonical operations (see, e.g., (Bodirsky et al., 2018; Bodirsky and Bodor, 2024; Mottet and Pinsker, 2020)), which is why we believe in the strong formulation of Conjecture 8.19. See Conjecture 9.3 for a conjecture that could hold for VCSPs in the more general setting of reducts of finitely bounded homogeneous structures.
@@
Let $C$ be the domain of $\Gamma$. Denote the relations $\operatorname{Opt}(R)$, $\operatorname{Opt}(S)$, $\operatorname{Opt}(T)$ by $R^{*}$, $S^{*}$, $T^{*}$, respectively. In the following, for $U\in\{R,S,T\}$ and variables $x,y$ we write $2U(x,y)$ for short instead of $U(x,y)+U(x,y)$. Let $\phi(a,b,c,d,e,f,g,h,i)$ be the $\tau$-expression
@@ 8.5. An example of formerly open complexity
There is a finitely bounded homogeneous dual ${\mathfrak{B}}$ of $\mu$ such that the valued $\tau$-structure $\Gamma:=\Gamma({\mathfrak{B}},\emptyset)$ has a binary fractional polymorphism which is canonical and pseudo cyclic with respect to $\operatorname{Aut}(\Gamma)$. Hence, $\operatorname{VCSP}(\Gamma)$ and the resilience problem for $\mu$ are in P. The polynomial-time tractability result even holds for resilience of $\mu$ with exogeneous relations from any $\sigma\subseteq\tau$.
@@
For every a finite $\tau$-structure ${\mathfrak{A}}$ that satisfies $\neg\mu$ and the sentences in (14) and (15), there exists a strong homomorphism to ${\mathfrak{B}}$, i.e., a homomorphism that also preserves the complements of $R$ and $S$. First observe that ${\mathfrak{B}}$ embeds the countably infinite complete graph, where $R$ is the edge relation and precisely one element lies in the relation $S$; this is because this structure maps homomorphically to ${\mathfrak{B}}$ and unless embedded, it contradicts ${\mathfrak{B}}\not\models\mu$. In particular, there are infinitely many $x\in B$ such that ${\mathfrak{B}}\models\neg S(x)\wedge\neg R(x,x)$ and by (15), for every $y\in B$, $x\neq y$, we have ${\mathfrak{B}}\models R(x,y)\wedge R(y,x)$.
@@
Every finite $\tau$-structure ${\mathfrak{A}}$ that satisfies $\neg\mu$ and the sentences in (14) and (15) embeds into ${\mathfrak{B}}$. In particular, ${\mathfrak{B}}$ is finitely bounded. Let ${\mathfrak{A}}$ be such a structure. By Theorem 8.14, there is an embedding $e$ of ${\mathfrak{A}}$ into ${\mathfrak{B}}_{\mu}$. Since ${\mathfrak{B}}_{\mu}$ is homogeneous and embeds every finite $\tau$-structure that satisfies $\neg\mu$, there exists a finite substructure ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{B}}_{\mu}$ satisfying the sentences in (14) and (15) such that $e({\mathfrak{A}})$ is a substructure of ${\mathfrak{A}}^{\prime}$ and for all distinct $a,b\in A$ there exists $s\in S^{{\mathfrak{A}}^{\prime}}$ such that ${\mathfrak{B}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. By Claim 1, there is a strong homomorphism $h$ from ${\mathfrak{A}}^{\prime}$ to ${\mathfrak{B}}$.
We claim that $h\circ e$ is injective and therefore an embedding of ${\mathfrak{A}}$ into ${\mathfrak{B}}$. Suppose there exist distinct $a,b\in A$ such that $h(e(a))=h(e(b))$. Since $e({\mathfrak{A}})$ satisfies the sentence in (14) and $h$ is a strong homomorphism, we obtain that ${\mathfrak{B}}_{\mu}\models R(e(a),e(a))\wedge R(e(b),e(b))$. Let $s\in S^{{\mathfrak{A}}^{\prime}}$ be such that ${\mathfrak{B}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. Hence,
@@
  • ${\mathfrak{M}},{\mathfrak{N}}\models\neg R((x_{1},y_{1}),(x_{2},y_{2}))$. Since ${\mathfrak{M}}$ and ${\mathfrak{N}}$ satisfy the sentences in (14) and (15) and ${\mathfrak{B}}$ satisfies (14) we must have ${\mathfrak{B}}\models\neg R(x_{1},x_{2})\wedge\neg R(y_{1},y_{2})$, and both sides of the inequality evaluate to $1$.
@@
It follows that $\omega$ is a binary fractional polymorphism of $\Gamma$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}(\Gamma)$. Polynomial-time tractability of $\operatorname{VCSP}(\Gamma)$ follows by Theorem 7.17 and 8.16. The final statement follows from Remark 8.17. ∎
@@ 9. Conclusion and Future Work
We formulated a general hardness condition for VCSPs of valued structures with an oligomorphic automorphism group and a new polynomial-time tractability result. We use the latter to resolve a resilience problem whose complexity was left open in the literature and conjecture that our conditions exactly capture the hard and easy resilience problems for conjunctive queries (under bag semantics), respectively. In fact, a full classification of resilience problems for conjunctive queries based on our approach seems feasible, but requires further research, as discussed in the following.
@@
One might hope to prove this conjecture under the assumption of the infinite-domain tractability conjecture. Recall that also the finite-domain VCSP classification was first proven conditionally on the finite-domain tractability conjecture (Kolmogorov et al., 2017; Kozik and Ochremiak, 2015), which was only confirmed later (Zhuk, 2017; Bulatov, 2017).
We also believe that the ‘meta-problem’ of deciding whether for a given conjunctive query the resilience problem is in P is decidable. This would follow from a positive answer to Conjecture 8.19 because $\Gamma^{*}_{m}$ can be computed and Item 4 of Proposition 7.16 for the finite-domain valued structure $\Gamma^{*}_{m}$ can be decided algorithmically using linear programming (Kolmogorov, 2019).
@@ Appendix A The Lebesgue Integral
Appendix B Regular Path Queries
We introduce (a Boolean version of) two-way regular path queries and show that they have finite duals. A two-way regular path query (2RPQ) over a finite relational signature $\tau$ is a regular expression over the alphabet $\Sigma_{\tau}$ that consists of the symbols $R$ and $R^{-}$ for every binary relation symbol $R\in\tau$. We assume that regular expressions are built up from alphabet symbols, $\emptyset$, $\varepsilon$, $+$ (union), ; (concatenation), and $\cdot^{*}$ (Kleene star). As usual, we may omit the concatenation symbol for brevity.
Every 2RPQ $\mu$ defines a regular language $L(\mu)$ over the alphabet $\Sigma_{\tau}$. It also defines a database query of arity 2, as follows. A path in a $\tau$-database $\mathfrak{A}$ is a sequence $a_{1}P_{1}a_{2}P_{2}\cdots P_{n-1}a_{n}$ with $a_{1},\dots,a_{n}\in A$ and $P_{1},\dots,P_{n-1}\in\Sigma_{\tau}$ such that for $1\leq i<n$, one of the following holds: (i) $P_{i}\in\tau$ and $\mathfrak{A}\models P_{i}(a_{i},a_{i+1})$ or (ii) $P_{i}=P^{-}$ with $P\in\tau$ and $\mathfrak{A}\models P(a_{i+1},a_{i})$. Note that we do not require $a_{1},\dots,a_{n}$ to be distinct. We call $a_{1}$ the source of the path, $a_{n}$ the target, and $P_{1}\cdots P_{n-1}$ the label. A pair $(a,b)\in A\times A$ is an answer to a 2RPQ $\mu$ if there is a path in $\mathfrak{A}$ that has source $a$, target $b$, and some label $w\in L(\mu)$. We use $\mu(\mathfrak{A})$ to denote the set of all answers to 2RPQ $\mu$ on database $\mathfrak{A}$ and write $\mathfrak{A}\models\mu$ if $\mu(\mathfrak{A})$ is non-empty. It is in this sense that our 2RPQs $\mu$ are Boolean: given a database $\mathfrak{A}$, we are interested in whether $\mathfrak{A}\models\mu$ instead of computing or verifying concrete answers.
Duals for 2RPQs are defined in exactly the same way as for conjunctive queries, that is, if $\mu$ is a 2RPQ with signature $\tau$, then a dual of $\mu$ is a $\tau$-structure ${\mathfrak{B}}$ such that any finite $\tau$-structure ${\mathfrak{A}}$ has a homomorphism to ${\mathfrak{B}}$ if and only if ${\mathfrak{A}}\not\models\mu$. We now argue that every 2RPQ has a finite dual. The proof refers to monadic Datalog programs, which we refrain from defining here in full detail. Note that the construction from (Calvanese et al., 2000) is related, but does not produce duals in our sense.
Lemma B.1.
Every 2RPQ has a finite dual.
Proof.
It was shown in (Bienvenu et al., 2014) that every simple Boolean monadic Datalog program has a dual. It thus suffices to show that every 2RPQ can be expressed as such a program.
$\begin{array}[]{rcl}P(x)&\leftarrow&R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{n}(\bar{x}_{n})\\[2.84526pt] P(x)&\leftarrow&\textsf{true}(x)\\[2.84526pt] \textsf{goal}()&\leftarrow&R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{n}(\bar{x}_{n})\end{array}$

### v4

@@ 8.5. The Resilience Tractability Conjecture
Let $\mu$ be a union of connected conjunctive queries over the signature $\tau$, and let $\sigma\subseteq\tau$. If the structure ($(\{0,1\};\operatorname{OIT})$ has no pp-construction in $\Gamma_{\mu,\sigma}$, then there exists a dual ${\mathfrak{B}}$ of $\mu$ such that $\operatorname{Aut}({\mathfrak{B}})$ contains an automorphism group of a finitely bounded homogeneous structure ${\mathfrak{A}}$ and $\Gamma:=\Gamma({\mathfrak{B}},\sigma)$ has a fractional polymorphism of arity $\ell\geq 2$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{A}})$ (and in this case, $\operatorname{VCSP}(\Gamma)$ is in P by Theorem 7.18).
[…]
Note that if ${\mathfrak{B}}$ is a relational structure on a finite domain $B$ and $\Gamma({\mathfrak{B}},\sigma)$ has a cyclic fractional polymorphism, then this polymorphism is trivially pseudo cyclic with respect to any group and it is canonical with respect to the trivial permutation group on $B$. Let $\tau=\{R_{b}\mid b\in B\}$ be a relational signature where all symbols are unary. Note that the relational structure ${\mathfrak{A}}=(B;(R_{b}^{{\mathfrak{A}}})_{b\in B})$ where $R_{b}^{{\mathfrak{A}}}=\{b\}$ for every $b\in B$ has a trivial automorphism group and is homogeneous. It is also finitely bounded: a finite relational $\tau$-structure embeds into ${\mathfrak{A}}$ if and only if it satisfies
$\forall x,y\left(\left(\bigvee_{b\in B}R_{b}(x)\right)\wedge\bigwedge_{b,c\in B,b\neq c}\big{(}\neg R_{b}(x)\vee\neg R_{c}(x)\big{)}\wedge\bigwedge_{b\in B}(R_{b}(x)\wedge R_{b}(y)\Rightarrow x=y)\right).$
Hence, every union of conjunctive queries each of which is homomorphically equivalent to a tree satisfies the condition from Conjecture 8.24 by Theorem 8.11 and Theorem 7.15.
Conjecture 8.24 is intentionally formulated only for VCSPs that stem from resilience problems, because it is known to be false for the more general situation of VCSPs for valued structures $\Gamma$ that have the same automorphisms as a reduct of a finitely bounded homogeneous structure (Bodirsky, 2021) (Section 12.9.1; the counterexample is even a CSP). However, the structures ${\mathfrak{B}}_{\mu}$ from Theorem 8.18 that allow to formulate resilience problems as VCSPs are particularly well-behaved for the universal-algebraic approach and more specifically, for canonical operations (see, e.g., (Bodirsky et al., 2018; Bodirsky and Bodor, 2024; Mottet and Pinsker, 2024)), which is why we believe in the strong formulation of Conjecture 8.24. See Conjecture 9.3 for a conjecture that could hold for VCSPs in the more general setting of reducts of finitely bounded homogeneous structures.
@@
Let $C$ be the domain of $\Gamma$. Denote the relations $\operatorname{Opt}(R)$, $\operatorname{Opt}(S)$, $\operatorname{Opt}(T)$ by $R^{*}$, $S^{*}$, $T^{*}$, respectively. In the following, for $U\in\{R,S,T\}$ and variables $x,y$ we write $2U(x,y)$ for short instead of $U(x,y)+U(x,y)$. Let $\phi(a,b,c,d,e,f,g,h,i)$ be the expression
@@ 8.6. An example of formerly open complexity
There is a finitely bounded homogeneous dual ${\mathfrak{B}}$ of $\mu$ such that the valued $\tau$-structure $\Gamma:=\Gamma({\mathfrak{B}},\emptyset)$ has a binary fractional polymorphism which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Hence, $\operatorname{VCSP}(\Gamma)$ and the resilience problem for $\mu$ are in P. The polynomial-time tractability result even holds for resilience of $\mu$ with exogeneous relations from any $\sigma\subseteq\tau$.
@@
For every finite $\tau$-structure ${\mathfrak{A}}$ that satisfies $\neg\mu$ and the sentences in (20) and (21), there exists a strong homomorphism to ${\mathfrak{B}}$, i.e., a homomorphism that also preserves the complements of $R$ and $S$. First observe that ${\mathfrak{B}}$ embeds the countably infinite complete graph, where $R$ is the edge relation and precisely one element lies in the relation $S$; this is because this structure maps homomorphically to ${\mathfrak{B}}$ and unless embedded, it contradicts ${\mathfrak{B}}\not\models\mu$. In particular, there are infinitely many $x\in B$ such that ${\mathfrak{B}}\models\neg S(x)\wedge\neg R(x,x)$ and by (21), for every $y\in B$, $x\neq y$, we have ${\mathfrak{B}}\models R(x,y)\wedge R(y,x)$.
@@
Every finite $\tau$-structure ${\mathfrak{A}}$ that satisfies $\neg\mu$ and the universal sentences in (20) and (21) embeds into ${\mathfrak{B}}$. In particular, ${\mathfrak{B}}$ is finitely bounded. Let ${\mathfrak{A}}$ be such a structure. By Theorem 8.19, there is an embedding $e$ of ${\mathfrak{A}}$ into ${\mathfrak{C}}_{\mu}$. Since ${\mathfrak{C}}_{\mu}$ is homogeneous and embeds every finite $\tau$-structure that satisfies $\neg\mu$, there exists a finite substructure ${\mathfrak{A}}^{\prime}$ of ${\mathfrak{C}}_{\mu}$ satisfying the sentences in (20) and (21) such that $e({\mathfrak{A}})$ is a substructure of ${\mathfrak{A}}^{\prime}$ and for all distinct $a,b\in A$ there exists $s\in S^{{\mathfrak{A}}^{\prime}}$ such that ${\mathfrak{C}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. By Claim 1, there is a strong homomorphism $h$ from ${\mathfrak{A}}^{\prime}$ to ${\mathfrak{B}}$.
We claim that $h\circ e$ is injective and therefore an embedding of ${\mathfrak{A}}$ into ${\mathfrak{B}}$. Suppose there exist distinct $a,b\in A$ such that $h(e(a))=h(e(b))$. Since $e({\mathfrak{A}})$ satisfies the sentence in (20), we have ${\mathfrak{B}}\models R(h(e(a)),h(e(a)))$. Let $s\in S^{{\mathfrak{A}}^{\prime}}$ be such that ${\mathfrak{C}}_{\mu}\models R(e(a),s)\wedge R(s,e(b))$. Hence,
@@
  • ${\mathfrak{M}},{\mathfrak{N}}\models\neg R((x_{1},y_{1}),(x_{2},y_{2}))$. We need to show that ${\mathfrak{B}}\models\neg R(x_{1},x_{2})\wedge\neg R(y_{1},y_{2})$. This is clear if $(x_{1},y_{1})=(x_{2},y_{2})$ by the definition of ${\mathfrak{N}}$. Suppose therefore that $(x_{1},y_{1})\neq(x_{2},y_{2})$. Since ${\mathfrak{M}}$ satisfies the sentence in (21), we have ${\mathfrak{M}}\models S(x_{1},y_{1})\wedge R((x_{2},y_{2}),(x_{2},y_{2}))$ or ${\mathfrak{M}}\models S(x_{2},y_{2})\wedge R((x_{1},y_{1}),(x_{1},y_{1}))$. Suppose that ${\mathfrak{M}}\models S(x_{1},y_{1})\wedge R((x_{2},y_{2}),(x_{2},y_{2}))$; the other case is analogous. Since ${\mathfrak{N}}$ satisfies the sentence in (21) as well, this implies that ${\mathfrak{N}}\models S(x_{1},y_{1})\wedge R((x_{2},y_{2}),(x_{2},y_{2}))$; note that if ${\mathfrak{N}}\models S(x_{2},y_{2})\wedge R((x_{1},y_{1}),(x_{1},y_{1}))$, we would get a contradiction with (19). By the definition of ${\mathfrak{M}}$ and ${\mathfrak{N}}$, we have ${\mathfrak{B}}\models S(x_{1})\wedge S(y_{1})\wedge R(x_{2},x_{2})\wedge R(y_{2},y_{2})$, in particular, by (19), $x_{1}\neq x_{2}$ and $y_{1}\neq y_{2}$. By (20), there is an $R$-edge in ${\mathfrak{B}}$ between $x_{1}$ and $x_{2}$ and between $y_{1}$ and $y_{2}$. By the condition (A) and (C) for ${\mathfrak{M}}$ and for ${\mathfrak{N}}$ with $1$ and $2$ switched, we see that ${\mathfrak{M}},{\mathfrak{N}}\models\neg R((x_{1},y_{1}),(x_{2},y_{2}))$ implies ${\mathfrak{B}}\models\neg R(x_{1},x_{2})\wedge\neg R(y_{1},y_{2})$. Therefore, both sides of the inequality evaluate to $1$.
@@
It follows that $\omega$ is a binary fractional polymorphism of $\Gamma$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Polynomial-time tractability of $\operatorname{VCSP}(\Gamma)$ follows by Theorem 7.18 and 8.20. The final statement follows from Remark 8.21. ∎
@@ 9. Conclusion and Future Work
We formulated a general hardness condition for VCSPs of valued structures with an oligomorphic automorphism group and a new polynomial-time tractability result. We use the latter to resolve the resilience problem for a conjunctive query whose complexity was left open in the literature and conjecture that our conditions exactly capture the hard and easy resilience problems for conjunctive queries (under bag semantics), respectively. In fact, a full classification of resilience problems for conjunctive queries based on our approach seems feasible, but requires further research, as discussed in the following.
@@
One might hope to prove this conjecture under the assumption of the infinite-domain tractability conjecture. Recall that also the finite-domain VCSP classification was first proven conditionally on the finite-domain tractability conjecture (Kolmogorov et al., 2017; Kozik and Ochremiak, 2015a), which was only confirmed later (Zhuk, 2017; Bulatov, 2017).
We also believe that the ‘meta-problem’ of deciding whether for a given conjunctive query the resilience problem is in P is decidable. This might follow from a positive answer to Conjecture 8.24 if an appropriate ${\mathfrak{B}}$ can be found effectively, because then $\Gamma^{*}_{{\mathfrak{B}},m}$ can be computed effectively and Item 4 of Proposition 7.17 for the finite-domain valued structure $\Gamma^{*}_{{\mathfrak{B}},m}$ can be decided using linear programming (Kolmogorov, 2019).